The foothaller kicked the ball and it was caught by the opposing goalkeeper, 4 metres above theground. The ball was travelling at 10 m/s Calculate the kinetic energy of the ball just before it was caught.


1
Expert's answer
2022-01-18T09:36:35-0500

The kinetic energy was


"KE=\dfrac 12mv^2=\dfrac 12·m·10^2=50m\text{ J}."

Substitute the mass to get the complete answer.